# Just Functioning as a Hook for Two-Stage Referring Multi-Object Tracking

### v2

@@ Abstract
Referring Multi-Object Tracking (RMOT) aims to localize target trajectories specified by natural language expressions in videos. Existing RMOT methods mainly follow two paradigms: one-stage strategies and two-stage ones. The former jointly trains tracking with referring but suffers from substantial computational overhead. Although the latter improves efficiency, it overlooks the inherent contextual aggregation capabilities of pre-trained visual backbones and takes a detour. Meanwhile, its fixed dual-tower architecture restricts compatibility with other visual / text backbones. To address these limitations, we propose JustHook, a novel hook-like framework for two-stage RMOT, which introduces two core components: (1) a Visual Feature Hook (VFH), enabling JustHook to extract context-rich local features directly from the original visual backbone like a hook; (2) a Parallel Combined Decoder (PCD), which transforms the passive cosine similarity measurement between independent modalities into active contrastive learning within the combined feature space. The proposed JustHook not only leverages the capabilities of pre-trained models but also breaks free from the constraints of inherent modality alignment, achieving strong scalability. Extensive experiments on Refer-KITTI and Refer-KITTI-V2 demonstrate that JustHook outperforms state-of-the-art methods across diverse encoder combinations, achieving a notable 7.77% HOTA improvement on Refer-KITTI-V2. Code will be made available soon.
1 Introduction
Unlike traditional cross-modal tasks such as image-text matching / retrieval [1; 2; 3], Referring Multi-Object Tracking (RMOT) [4] establishes dynamic correspondence between textual descriptions and spatiotemporal trajectories of one or multiple unique targets in video sequences. In RMOT, the tracker is required not only to establish a correspondence between the expressions and the objects, but also to track these objects throughout the video. For instance, given an expression like “black car on the left", the model is required to provide the trajectories of all black cars on the left side of the video while filtering out semantically irrelevant objects. The introduction of natural language makes RMOT require both a fine-grained understanding of the specific attributes (like color) and the grasp of targets’ contextual information (like location) within the entire video.
Figure 1: The comparison between previous two-stage RMOT frameworks and ours. (a) illustrates existing two-stage methods, which utilize the CLIP visual-textual feature space and feed both global and cropped images into the image encoder to enhance the contextual information. (b) presents the proposed JustHook, which operates like a hook to directly extract context-rich local features from any visual backbone without the need for additional modules or reusing the backbone.
Existing RMOT methods mainly follow two architectural paradigms: one-stage and two-stage. Current one-stage methods [4; 5; 6] connect the tracker and the matcher in series through soft decoupling, which introduce auxiliary classification modules after the conventional MOT architecture MOTR [7] to verify the correspondence between expressions and targets. Although they enable end-to-end referring tracking, the relatively high computational overhead of MOTR prevents their widespread application in practical scenarios. In contrast, two-stage methods [8] adopt hard decoupling by developing specialized cross-modal matchers that operate on trajectory proposals generated by off-the-shelf trackers. This modular design inherently provides greater flexibility in tracker selection and scalability across diverse scenarios.
However, in their pursuit of architectural decoupling from trackers, current representative two-stage methods, taking iKUN [8] as an example, actually make a detour and lack the full utilization of visual backbones. As shown in Fig. 1 (a), to meet the requirements of RMOT, iKUN reuses the CLIP image encoder to encode the global image and the cropped target image. Then, an additional local-global interaction module is introduced to extract the context-rich target visual features. Finally, to correct the semantic space offset generated during this process, it also needs to further modulate the visual features using the text features. Therefore, it inevitably damages the carefully constructed semantic space of CLIP, thus resulting in iKUN’s inability to generalize to the complex Refer-KITTI-V2 dataset (with only 10.32% HOTA and 2.17% DetA). In fact, most visual backbones (e.g., ResNet [9], Swin [10]) exhibit excellent contextual capabilities inherently through progressive (rather than abrupt) receptive field expansion. Even CLIP [1], which is pre-trained at the image level, still possesses fine-grained capabilities [11; 12].
In addition, to easily leverage the capabilities of the pre-trained CLIP, existing two-stage frameworks are directly designed based on the dual-tower structure of CLIP. As illustrated in Fig. 1 (a), their passive approach of measuring alignment between visual and text modalities via cosine similarity confines current two-stage RMOT methods to CLIP’s semantic space, while CLIP itself inherently possesses modality alignment capabilities. Such a modality-independent perspective greatly limits the integration with other advanced encoders and fails to achieve future-proofing.
To address these issues, we propose a novel hook-like framework called JustHook as shown in Fig. 1 (b). The proposed Visual Feature Hook (VFH) operates like a hook, fully preserving the original workflow of visual backbones while extracting context-rich local features without reusing the backbone. Its efficient extraction capability and fully decoupled design enable JustHook to fully unleash the performance of the pre-trained model and integrate seamlessly with various off-the-shelf visual backbones. In VFH, we design an Optical-flow-based Motion Extractor (OME) to replace iKUN’s simplistic temporal pooling to obtain more accurate motion information. Then, we further transform the previous dual-tower structure into an encoder-decoder structure. Using the proposed Parallel Combined Decoder (PCD), JustHook converts the original passive measurement between independent modalities into an active inspection within the combined feature space, making the two-stage RMOT no longer strongly dependent on the inherent semantic space of CLIP. Ultimately, the proposed JustHook framework enables coupling with any visual backbone and any text encoder, and outperforms state-of-the-art (SOTA) methods on widely used Refer-KITTI [4] and Refer-KITTI-v2 [5] datasets. Our contributions can be summarized as follows:
(1) We propose a novel hook-like framework called JustHook, which directly extracts context-rich local features from the original encoding process of the visual backbone without additional modules or reusing the backbone, thus more fully unleashing the knowledge in pre-trained models.
(2) We transform current two-stage RMOT methods from modality-separated measurement into understanding and classifying sample pairs within a combined space, which eliminates the dependency on the inherent modality alignment of CLIP, thereby enhancing the performance and adaptability to new advancements.
(3) By efficiently extracting visual features and decoding in the combined feature space, the proposed VFH and PCD modules achieve effective coupling with arbitrary encoders and obtain robust performance improvement.
(4) JustHook outperforms state-of-the-art (SOTA) methods in both performance and efficiency on widely used RMOT datasets Refer-KITTI and Refer-KITTI-V2. It achieves a 7.77% improvement in HOTA metric and performs well across various combinations of visual / text encoders, demonstrating its superiority and robustness.
2 Related Work
2.1 Referring Single Object Tracking
Referring Single Object Tracking (RSOT) has been studied for several years. Most SOTA methods [13; 14; 15] follow classic tracking strategies, integrating advancements in referring video segmentation [16] (or detection [17]) and focusing on end-to-end joint tracking algorithms. JointVLT [18] guides the model to perform grounding in the first frame by inputting only a zero-pad, and then continuously tracks the target in subsequent frames based on template and expression. TransVLT [19] uses a set of proxy tokens for cross-modal information fusion, which project linguistic features and generate kernel vectors to modulate visual features. Instead of directly decoding the bounding boxes, MMTrack [20] constructs a set of queries from expression and bounding box cues and generates the target trajectories using an autoregressive approach. Although RSOT has made significant achievements [21; 22], its development is limited by highly limited task scenarios.
2.2 Referring Multi Object Tracking
TransRMOT [4] has pioneered RMOT by introducing the Refer-KITTI dataset and a baseline model based on MOTR [7]. Building on TransRMOT [4], MLS-Track [6] further designs a multi-level semantic interaction module to enhance the interaction between semantic and visual information. iKUN [8] decouples tracker and matcher, developing a modular framework that achieves superior performance compared to end-to-end baselines. Based upon recent advances, TempRMOT [5] constructs the enhanced Refer-KITTI-v2 dataset by leveraging large language models [23], achieving SOTA performance, while its computational overhead still persists. While iKUN imposes a lower training burden, its overly straightforward approach fails to effectively harness the capabilities of pre-trained models. Moreover, iKUN [8] confines existing two-stage frameworks within the CLIP [1] semantic space, rendering them non-future-proof.
3 Method
@@
3.1 Problem Formulation
As a two-stage method, JustHook is typically integrated directly after any off-the-shelf tracker. Generally, we use a track-by-detection (TBD) tracker to maximize the efficiency advantages. Each tracked target will be matched with multiple expressions by JustHook. For a target in the current frame $I_{t}$, JustHook utilizes its trajectory over the previous $T-1$ frames, forming an ordered sequence of $T$ frames: $\{I_{t-T+1},\dots,I_{t-1},I_{t}\}\in\mathbb{R}^{T\times 3\times H\times W}$. Besides images, JustHook also takes the sequence of bounding box coordinates for the target in each frame as input: $\{B_{t-T+1},\dots,B_{t-1},B_{t}\}\in\mathbb{R}^{T\times 4}$, which are directly derived from the tracker. At each step, a set of $N$ expressions is sampled, and their linguistic features are extracted from the text encoder: $\{E_{1},E_{2},\dots,E_{N}\}\in\mathbb{R}^{N\times L\times C_{t}}$, where $L$ represents the length of the expression. Finally, JustHook outputs a set of logits $O\in\mathbb{R}^{N\times 2}$, where each row corresponds to the referring result between one of the $N$ expressions and the current target.
3.2 A Hook-liked Framework
Hook is a mechanism in a program or system that allows for intercepting, monitoring, or modifying specific events or operations by registering callback functions to intervene in and extend certain processes [24]. As shown in Fig. 2, JustHook solves the two-stage RMOT problem by working like a Hook. Besides off-the-shelf visual backbones and text encoders, it is composed of only two parts, namely the Visual Feature Hook (VFH) and the Parallel Combined Decoder (PCD), which is simple yet effective. In the VFH, JustHook constructs a target position grid using bounding boxes and extracts target features from multi-scale visual feature maps through bilinear grid sampling. On the one hand, direct feature sampling enables VFH to efficiently extract local features that incorporate contextual information aggregated by the visual backbone. On the other hand, since bilinear grid sampling does not impede gradient propagation, VFH can leverage the gradient pathways of the backbone network to adaptively adjust the contextual information of local features. Subsequently, the Optim-Flow-Based Motion Extractor (OME) aggregates the local features from multiple frames into trajectory features. Finally, the PCD integrates bimodal features and directly decodes pair-wise matching scores within the combined feature space. Simultaneously, it processes each target against multiple texts in parallel, which not only boosts inference efficiency but also elevates performance via contrastive learning.
3.3 Visual Feature Hook
Figure 3: Illustration of bilinear grid sample. Leveraging the four nearest feature pixels and their coordinates, grid sampling can compute features at arbitrary continuous coordinates via bilinear interpolation.
@@ 3.3 Visual Feature Hook
Context-Rich Local Features
As shown in Fig. 4, in VFH, we first convert the input bounding box sequence $B=<x_{1},y_{1},w_{b},h_{b}>\in\mathbb{R}^{4}$ into a continuous coordinate grid $P\in\mathbb{R}^{2\times h_{i}\times w_{i}}$, where $(x_{1},y_{1})$ represents the coordinates of the top-left corner of the target, $w_{b}$ and $h_{b}$ are the width and height of the bounding box, and $w_{i}$ and $h_{i}$ represent the width and height of the target features in the $i$-th layer. The elements $P_{i,j}$ in the coordinate grid are calculated as follows:
$P_{i,j}$ $=P_{1,1}+(i-1)\cdot\Delta x+(j-1)\cdot\Delta y$ (1)
$\Delta x$ $=(\frac{w_{b}}{w_{i}},0),\Delta y=(0,\frac{h_{b}}{h_{i}})$
where, $P_{1,1}=(x_{1},y_{1})$, and $\Delta x$ and $\Delta y$ represent the stride between sampling points. Taking the sampled pixel $X\in\mathbb{R}^{C}_{i}$ as an example, its coordinate is $P_{x}=(\hat{x},\hat{y})\in\mathbb{R}^{2}$. Through the four feature pixels $A$, $B$, $C$, $D$ on the original feature map that are closest to $X$, $X$ can be calculated using the following bilinear interpolation:
$X=(1-w)(1-h)A+w(1-h)B+(1-w)hC+whD$ (2)
where the coordinate of the lower-left pixel $A$ is $P_{A}=(x,y)\in\mathbb{N}^{2}$. Correspondingly, the coordinate of the upper-right point $D$ is $P_{A}=(x+1,y+1)\in\mathbb{N}^{2}$. And $w=\hat{x}-x$, $h=\hat{y}-y$ are respectively the offset of $X$ in the horizontal and vertical directions relative to the point $A$.
Direct feature-level sampling allows VFH to efficiently harness the intrinsic contextual capabilities of pre-trained models. Meanwhile, by preserving gradient flow continuity, the grid sample ensures that although VFH only samples features from the target region, gradients can still propagate globally via the backbone network. This enables VFH to inherently acquire the contextual knowledge that has been integrated by pre-trained networks and adjust it simply by encoding the global image. It becomes particularly pronounced in dense visual backbones as well as Transformer-based backbones designed for long-range modeling.
Optim-flow-based motion extractor
Motion information is an indispensable part of processing motion expressions (e.g., "the car turning left") in RMOT. However, the existing two-stage methods only use simple temporal pooling to compress multi-frame information, ignoring the target optical flow naturally provided by the trajectory proposals in the two-stage RMOT, which is one of the most widely used and effective methods for describing the motion information in a visual scene. Therefore, we concatenate along the channel dimension both the coordinate grid displacements $\Delta Grid=Cat(\{P_{t-i}-P_{t-i-1}\}_{i=0}^{T})$, which serve as explicit optical flow (the blue part in Fig. 4), and the multi-frame Context-Rich Local Features, which serve as feature optical flow (the purple part in Fig. 4), thereby effectively utilizing the trajectory proposals to form target optical flow features. Meanwhile, to further enhance the model’s perception capability in the spatial dimension, we design three different strategies: (1) MLP strategy, where features are simply fused using a multi-Layer perceptron (MLP); (2) Conv strategy, which applies convolution after channel-wise concatenation to focus on the local information of the target; and (3) Attention strategy, which leverages inter-frame self-attention [25] to adaptively highlight regions of interest.
3.4 Parallel Combined Decoder
Figure 5: Illustration of the PCD. Compared with the passive strategy of cosine similarity for single sample pairs in (a), PCD actively examines the matching situations of combined sample pairs in the manner shown in (b) and conducts contrastive learning, no longer relying on the modality alignment capability inherent in CLIP. In (c), the composition of PCD is specifically introduced, which is mainly composed of mask cross-attention and FFN.
To mitigate the dependency on the modality alignment of encoders and keep future-proofing, we adopt an encoder-decoder strategy within the combined space instead of the cosine similarity-based score computation between independent modalities. As illustrated in Fig. 5 (b), the proposed Parallel Combined Decoder (PCD) operates on unified semantic representations formed by the concatenation of visual features and expression words. Rather than treating each modality separately, the PCD takes the combination of the two modalities as the basic feature space. In this way, the cross-modal matching is transformed into a task of understanding and classifying sample pairs within a combined space.
@@ 3.4 Parallel Combined Decoder
Specifically, as shown in Fig. 5 (c), after concatenating the trajectory features $J\in\mathbb{R}^{C_{t}\times h_{i}\times w_{i}}$ and linguistic features $E\in\mathbb{R}^{N\times L\times C_{t}}$ in the spatial dimension (the word dimension in $E$), PCD uses them as Key and Value, with the learnable query vectors $Q\in\mathbb{R}^{N\times C_{t}}$ as the Query, to perform masked cross-attention. The extracted information is then further summarized through a Feed-Forward Network (FFN) followed by an MLP to generate the match scores $O\in\mathbb{R}^{N\times 2}$. In the cross-attention mask $A\in\mathbb{R}^{N\times(h_{i}\cdot w_{i}+N\cdot L)}$, we assign expressions to the query vectors one by one and ensure that the visual semantics are always visible to all query vectors. At each layer, we iteratively decode queries from shallow to deep, facilitating the analysis of the target state at different feature granularity.
Through the delicate design of the attention mask, PCD further extends the previous serial processing in Fig. 5 (a) into a parallel processing paradigm for multiple expressions, which enables explicit contrastive learning between sample pairs and improves the inference efficiency. Meanwhile, compared with relying on the inherent classification hyperplane of the modality-aligned semantic spaces, the approach of actively solving the hyperplane by combining bimodal features is more suitable for pairing with various emerging encoders, liberating two-stage RMOT from the dependence on CLIP.
@@
Table 1: Comparison with state-of-the-art methods on Refer-KITTI [4] and Refer-KITTI-V2 [5].
@@
Two-Stage FairMOT [28] 22.78 14.43 39.11 74.77
DeepSORT [29] 25.59 19.76 34.31 71.34
ByteTrack [30] 24.95 15.50 43.11 73.90
@@
JustHook CLIP [1] CLIP-R 50 [1] 52.93 42.36 66.22 90.82
BERT [33] Swin-T [10] 51.12 39.85 65.75 89.58
Resnet34 [9] 53.15 42.77 66.23 89.72
ROPE Swin-T [34] 53.30 42.44 67.13 89.46
RoBERTa [35] Swin-T [10] 51.03 39.50 65.99 90.98
Resnet34 [9] 53.28 41.89 67.87 90.75
ROPE Swin-T [34] 53.83 43.40 66.86 90.85
@@
JustHook CLIP [1] CLIP-R 50 [1] 41.44 28.71 59.95 89.12
BERT [33] Swin-T [10] 41.14 28.78 58.94 88.97
Resnet34 [9] 41.04 29.24 57.76 89.19
ROPE Swin-T [34] 41.16 29.41 57.76 89.52
RoBERTa [35] Swin-T [10] 40.32 27.47 59.29 88.96
Resnet34 [9] 41.71 30.07 57.99 89.25
ROPE Swin-T [34] 42.81 31.35 58.61 89.22
3.5 Loss Computation

### v3

@@ Abstract
Referring Multi-Object Tracking (RMOT) aims to localize target trajectories in videos specified by natural language expressions. Despite recent progress, the intrinsic relationship between the two subtasks of tracking and referring in RMOT has not been fully studied. In this paper, we present a systematic analysis of their interdependence, revealing that current two-stage Referring-by-Tracking (RBT) frameworks remain fundamentally limited by insufficient modeling of subtask interactions and inflexible reliance on semantic alignment modules like CLIP. To this end, we propose JustHook, a novel two-stage RBT framework where a Hook module is firstly designed to redefine the linkage between subtasks. The Hook is built centered on grid sampling at the feature-level and is used for context-aware target feature extraction. Moreover, we propose a Parallel Combined Decoder (PCD) that learns in a unified joint feature space rather than relying on pre-defined cross-modal embeddings. Our design not only enhances the interpretability and modularity but also significantly improves the generalization. Extensive experiments on Refer-KITTI, Refer-KITTI-V2, and Refer-Dance demonstrate that JustHook achieves state-of-the-art performance, improving the HOTA by +6.9% on Refer-KITTI-V2 with superior efficiency. Code will be available soon.
Introduction
By integrating natural language processing into traditional Multi-Object Tracking (MOT), Referring Multi-Object Tracking (RMOT) [33] aims to achieve consistent tracking of specific targets through precise guidance provided by natural language. Several recently proposed datasets and studies have revealed the progress. However, a deeply systematic task analysis is still lacking to support the development of a generalizable methodology. In this paper, we delve into a critical yet overlooked issue in RMOT: As a typical cross-task problem, how should the tracking task and referring task be effectively associated?
Figure 1: (a) Referring-by-Tracking classifies MOTR [35]-decoded queries via a linear layer. (b) Tracking-by-Referring associates trajectories within boxes located by GroundingDINO [20].
Although rarely discussed explicitly, as shown in Fig. 1, the underlying logic between subtasks has always been a central focus of research in RMOT. First introduced by TransRMOT [33] and extended by TempRMOT [39], the one-stage Referring-by-Tracking (RBT) paradigms pioneer an end-to-end design that prioritizes the tracking task, where matching scores are computed directly using trajectory queries decoded by MOTR [35]. In contrast, LaMOTer [17], highlights the referring task and constructs trajectories by linking detection boxes predicted by GroundingDINO [20] (i.e. Tracking-by-Referring, TBR). Despite being designed for different scenarios (TempRMOT for closed-set and LaMOTer for open-set), both methods achieve impressive performance within their respective domains.
Figure 2: Comparison of subtask logic in existing methods. (a) previous two-stage RBT reuses the encoder to joint subtask and computes CLIP-based cosine similarity as matching score. (b) Our method directly hooks features from the visual backbone and decodes scores via PCD in a composed feature space.
However, both MOTR [35] and GroundingDINO [17] often suffer from significant computational overhead. Among existing methods, the two-stage RBT paradigm, exemplified by iKUN [7], addresses efficiency by first employing lightweight off-the-shelf trackers to generate trajectory candidates, followed by an independent referring stage as shown in Fig. 2 (a). Nevertheless, this approach has gradually lost its popularity due to its poor generalization performance on complex datasets.
Rethinking these paradigms from the perspective of inter-subtask, we observe that both the popular one-stage RBT paradigm and the TBR paradigm, despite their differing logics, adhere to a common principle: ensuring the sufficiency and necessity of information transmission between preceding and succeeding tasks. Sufficiency implies that the subsequent task can receive all the information it needs, while necessity ensures that the preceding task transmits only the necessary information, without redundancy. In one-stage RBT, the end-to-end design provides a smooth gradient flow, allowing the trajectory queries to adaptively adjust the transmitted information according to downstream demands. In TBR, although lacking continuous gradient propagation, it reduces redundancy by placing the tracking task, which naturally requires less information, after the referring stage. In contrast, current two-stage RBT approaches, despite improving efficiency, only achieve suboptimal sufficiency and low necessity. Unlike the inherent advantages of other paradigms, current two-stage RBT methods can neither offer seamless information flow between subtasks nor eliminate the need to prioritize a lightweight tracker for efficiency. Therefore, during bridging, given that only trajectory boxes and raw images are available as prior information, and meanwhile, the referring subtask involves both local attributes (e.g., color, category) and global spatial cues (e.g., position, direction), these methods have to reuse the encoder to process both full images and cropped target patches via CLIP [28] as illustrated in Fig. 2(a), thereby focusing on the specific target while enriching the spatial context.
However, a widely known fact but neglected by them is that most visual backbones (e.g., ResNet [12], Swin [22]) inherently possess strong spatial contextual modeling capabilities. Even CLIP, though pre-trained at the image-level, has demonstrated fine-grained recognition abilities [31; 16]. Hence, neglecting the pre-existing capabilities and instead introducing additional re-training modules is a suboptimal approach to achieving sufficient information propagation. The irrelevant content in global images and the linking of subtasks by reusing the encoder also undermine the necessity. The 38.52% drop in HOTA of iKUN [7] on the Refer-KITTI-v2 [39] benchmark further demonstrates that unnatural information flow fails to fully exploit the backbone’s potential and compromises its robustness.
Therefore, rethinking the coupling between subtasks in a more deliberate and holistic perspective, a better bridging process should (1) precisely extract target regions by using tracking boxes to maximize necessity, and (2) preserve backbone’s spatial gradient flow to obtain contextual information and achieve better sufficiency. Based on these two principles, we find that nothing fits them better than the grid sampling operation. Therefore, centered around it, we build a new two-stage RBT framework, named JustHook, whose name draws inspiration from hook functions in programming [23]. Although grid sampling has been widely used in other methods such as ROI pooling [29] and Deformable Attention [43], its role in RMOT remains unexplored. To the best of our knowledge, we are the first to identify its suitability for RMOT and unlock its potential through carefully designed supporting modules.
Furthermore, to leverage the advantages of the new bridging module on dense visual backbones, we propose an active combined feature metric to replace the traditional cosine similarity matching by introducing the Parallel Combined Decoder (PCD). Finally, JustHook can be coupled with any backbone and text encoder and achieve robust performance. Our contributions can be summarized as follows:
(1) We offer a new perspective—subtask correlations, to reevaluate existing RMOT paradigms and reveal the inherent flaws in current two-stage RBT frameworks.
(2) Rethinking the subtask correlation patterns in the two-stage RBT paradigm, we propose a new framework, JustHook, which establishes well-suited information pathways and enables flexible coupling with arbitrary encoders.
(3) JustHook outperforms SOTA methods on the Refer-Dance [7], Refer-KITTI [33] and Refer-KITTI-V2 [39] datasets. It achieves a notable 6.9% HOTA improvement on Refer-KITTI-V2 and performs the highest efficiency.
Related Work
@@ Introduction
Referring Single Object Tracking
Referring Single Object Tracking (RSOT) has been studied for several years. Most SOTA methods [2; 34; 10; 9] follow classic tracking strategies, integrating advancements in referring video segmentation [6] (or detection [32]) and focusing on end-to-end joint tracking algorithms. JointVLT [40] performs grounding in the first frame by inputting only a zero-pad, and then continuously tracks the target in subsequent frames based on template and expression. TransVLT [42] uses a set of proxy tokens for cross-modal information fusion, which project linguistic features and generate kernel vectors to modulate visual features. Instead of directly decoding the bounding boxes, MMTrack [41] constructs a set of queries from expression and bounding box cues and generates the target trajectories using an auto-regressive approach. Although significant achievements [15; 36] made, the development of RSOT is limited by highly limited task scenarios.
Referring Multi Object Tracking
TransRMOT [33] pioneered RMOT by introducing the Refer-KITTI dataset along with a baseline model based on MOTR [35]. Building upon TransRMOT, MLS-Track [26] further enhances cross-modal interaction through a multi-level module design. To address the high computational cost of these methods, iKUN [7] innovatively proposed a modular two-stage RBT paradigm. Recently, TempRMOT [39] constructs the enhanced Refer-KITTI-V2 dataset by leveraging large language models [3]. Focusing on RMOT performance under open-vocabulary scenarios, LaMOTer introduces a Tracking-by-Referring architecture to further enrich the subtask logic of RMOT. While each paradigm has its own strengths and weaknesses, few studies have examined the internal interaction between subtasks in RMOT or identified the limitations of current two-stage RBT methods.
Method
Problem Formulation
In RMOT, the trajectory of target $i$ can be represented by its sequence of detection boxes across frames, denoted $\mathcal{B}^{i}=\{B^{i}_{t}\}_{t=t^{i}}^{T^{i}}$, where $t$ indicates the frame index, and $t^{i}$ and $T^{i}$ denote the starting and ending frames in which the target appears. We use $\mathcal{B}^{i}_{x:y}$ to denote a segment of target $i$’s trajectory, where $x$ and $y$ represent the starting and ending frames of this segment, respectively. The index $t$ is not necessarily continuous, as the target may disappear or be occluded in certain frames and reappear later. For each video, let $\mathcal{O}=\{\mathcal{B}_{i}\}_{i=1}^{K}$ be the set of all $K$ target trajectories, and let $\mathcal{E}=\{E_{j}\}_{j=1}^{N}$ denote the set of $N$ natural language expressions annotated for the targets in the whole video. The process of inferring all target trajectories $\mathcal{O}$ from the video frame sequence $\mathcal{I}=\{I_{t}\}_{t=0}^{T}$ constitutes the tracking subtask, while the referring subtask can be formulated as learning a many-to-many matching relation between the expression set $\mathcal{E}$ and the trajectory set $\mathcal{O}$:
$\mathcal{R}\subseteq\mathcal{E}\times\mathcal{O}$
where $(E_{j},\mathcal{B}^{i}_{x:y})\in\mathcal{R}$ indicates that expression $E_{j}$ refers to trajectory segment $\mathcal{B}^{i}_{x:y}$. Noticed that the reason why we use a segment instead of the entire trajectory is that certain expressions, such as “people on the left”, may become invalid as the target moves. In the two-stage RBT paradigm, the tracking subtask is handled by an off-the-shelf tracker, and we focus solely on the bridging module and the referring subtask.
Overview
To facilitate understanding, as illustrated in Fig. 3, we decompose JustHook into three main components: (1) Hook, a coupling module centered around grid sampling; (2) OME, an Optical-flow-based Motion Extractor; and (3) PCD, a Parallel Combined Decoder. Given a $p$-frame segment of a trajectory, denoted $\hat{\mathcal{B}}_{t:t+p}$, we sample $\hat{N}$ expressions $\hat{\mathcal{E}}=\{E_{j}\}_{j=1}^{\hat{N}}$ from the expression set $\mathcal{E}$ during training, while during inference, we iterate over all expressions. JustHook ultimately outputs $\hat{N}$ matching scores per forward, each corresponding to the match between the trajectory segment $\hat{\mathcal{B}}_{t:t+p}$ and one of the expressions in $\hat{\mathcal{E}}$.
Hook
Serving as a bridging module, Hook is designed to receive the output from the tracking subtask, denoted as $\hat{\mathcal{B}}_{t:t+p}$, and extract features required by the referring subtask. Unlike previous approaches that rely on additional modules or encoding procedures, Hook directly grid-samples the target features from the raw encoding stream of the backbone at the feature level. Its smooth gradient path makes full use of the contextual aggregation capability of pretrained models. Hook consists of two components: a neighboring grid sampling specifically adapted for RMOT, and a long-range modeling enhancement module.
Neighboring Grid Sample.
Naive grid sampling typically relies on hand-crafted or learnable coordinate grids. In RMOT, however, the coordinate grid can be naturally obtained from the tracking results $\hat{\mathcal{B}}_{t:t+p}$. Given a coordinate grid $P$, the value at any point $X$ is obtained by interpolating its four surrounding feature values, denoted $Y_{a},Y_{b},Y_{c},$ and $Y_{d}$. The interpolation weights $D_{a},D_{b},D_{c},$ and $D_{d}$ depend on the offsets, and satisfy $D_{a}+D_{b}+D_{c}+D_{d}=1$. Taking $Y_{a}$, the top-left neighbor of $X$, as an example: if its normalized horizontal and vertical offsets relative to $X$ are $(u,v)$, then its interpolation weight is $D_{a}=(1-u)(1-v)$. $X$ can be calculated as follows:
$X=D_{a}\cdot Y_{a}+D_{b}\cdot Y_{b}+D_{c}\cdot Y_{c}+D_{d}\cdot Y_{d}$
While naive grid sampling is already well-aligned with the requirements of the bridging module, a notable distributional gap exists in the two-stage RBT paradigm: the referring subtask is typically trained with ground-truth trajectories, whereas inference relies on trajectory candidates produced by an off-the-shelf tracker. To narrow this gap, we perturb the coordinate grids to simulate potential tracking errors. Specifically, we adopt three operations (i) randomly sample fragmented trajectory segments to emulate target loss, (ii) inject Gaussian noise to model localization inaccuracies, and (iii) re-combine grid sequences within a batch to mimic identity switches. This augmented strategy enables the model to acquire neighborhood awareness across temporal, spatial, and instance dimensions, thereby reducing overfitting. Furthermore, spatial neighbor sampling enhances the model’s capacity to capture broader contextual information.
Figure 4: Illustration of Long-Range Enhancement. Guided by the linguistic feature $F_{l}$, we computes $M$ reference points $P_{r}$ through a Transformer decoder and a residual fully connected network, which then are used to actively provide reference features by grid sampling.
Long-Range Enhancement.
Although grid sampling ensures a smooth gradient path, the model’s perception of external features will weaken as the spatial distance from the sampled region increases. To address this, we design an additional long-range modeling enhancement module for Hook to help the model actively acquire global features.
As illustrated in Fig. 4, we incorporate linguistic priors by leveraging expression information and perform cross-attention between learnable query vectors $Q_{LR}\in\mathbb{R}^{\hat{N}\times M\times C}$ and the linguistic features $F_{l}\in\mathbb{R}^{\hat{N}\times L\times C}$ output by the text encoder. The resulting attention output is then passed through an multi-Layer perceptron (MLP) and a sigmoid layer to generate normalized 2D reference points $P_{r}$. These reference points are used alongside the coordinate grid $P$ to extract visual features via grid sampling. The sampled features are then temporally pooled to form the reference feature $F_{r}\in\mathbb{R}^{\hat{N}\times M\times C}$. Finally, in PCD, the reference features are fused with the combined features to compute the matching score.
Optical-flow-based Motion Extractor
Motion information is an indispensable part of processing motion expressions (e.g., “the car turning left”) in RMOT. However, the existing methods only use simple temporal pooling to compress multi-frame information [7], ignoring the target optical flow naturally provided by the trajectory proposals. Therefore, we concatenate along the channel dimension both the coordinate grid displacements $\Delta Grid=Cat(\{P_{t-i}-P_{t-i-1}\}_{i=0}^{T})$, which serve as explicit optical flow (the blue part in Fig. 5), and the multi-frame features $F_{v}\in\mathbb{R}^{p\times C\times H\times W}$, which serve as feature optical flow (the purple part in Fig. 5). To further enhance the model’s perception capability in the spatial dimension, we design three different strategies: (1) MLP strategy, where features are simply fused using a multi-Layer perceptron (MLP); (2) Conv strategy, which applies convolution after channel-wise concatenation to focus on the local information of the target; and (3) Attention strategy, which leverages inter-frame self-attention [30] to adaptively highlight regions of interest.
@@
Parallel Combined Decoder
The current two-stage RBT paradigm usually takes CLIP [28] as a mandatory component, using its inherent cross-modal alignment capability to directly measure the matching between trajectory features $F_{j}$ and linguistic features $F_{l}$ via cosine similarity. On one hand, the theoretical upper bound of CLIP limits the potential of previous methods. On the other hand, our grid-sampling-based hook tends to perform better with backbones that excel in dense and long-range modeling scenarios. Therefore, as illustrated in Fig. 6 (b), we introduce a Parallel Combined Decoder (PCD) that operates on a joint embedding space formed by concatenating visual features and expression words, allowing the model to leverage the inherent differences between positive and negative pairs instead of relying on hard-to-measure cross-modal similarity in a misaligned space.
@@
Specifically, as shown in Fig. 6, PCD concatenates the trajectory features $F_{j}\in\mathbb{R}^{C\times H\times W}$ and linguistic features $F_{l}\in\mathbb{R}^{\hat{N}\times L\times C}$ in the spatial dimension (the word dimension in $F_{l}$), and then uses them as the Key and Value, with the learnable query vectors $Q\in\mathbb{R}^{\hat{N}\times C}$ as the Query, to perform masked cross-attention. Under the long-range enhancement setting, we further include the reference features $F_{r}\in\mathbb{R}^{\hat{N}\times M\times C}$ in the concatenation to actively extract global information. The decoded queries are then further summarized through a Feed-Forward Network (FFN) followed by an MLP to generate the match scores $O\in\mathbb{R}^{\hat{N}\times 2}$.We control the cross-attention mask $A\in\mathbb{R}^{\hat{N}\times(H\cdot W+\hat{N}\cdot L+\hat{N}\cdot M)}$ so that each query vector can attend only to its corresponding textual features and reference features, as well as all visual features.
@@
Table 1: Comparison with state-of-the-art methods on Refer-KITTI [33] and Refer-KITTI-V2 [39]. “” indicates that long-range enhancement is not used.
@@
Two-Stage ByteTrack [37] 24.95 15.50 43.11 73.90
@@
JustHook CLIP [28] CLIP-R 50 [28] 53.49 43.62 65.68 90.90
RoBERTa [21] Resnet34 [12] 53.47 43.01 66.55 90.95
ROPE Swin-T [14] 53.91 44.02 66.09 90.97
@@
JustHook CLIP [28] CLIP-R 50 [28] 41.03 29.08 58.02 89.03
RoBERTa [21] Resnet34 [12] 41.50 29.21 59.10 89.25
ROPE Swin-T [14] 41.94 30.45 57.92 89.00
Figure 6: Illustration of the PCD. Compared with the passive strategy of cosine similarity, PCD actively examines the matching of combined sample pairs by concatenation and conducts contrastive learning, no longer relying on the modality alignment capability inherent in CLIP.
Loss Computation